Advertisements
Advertisements
प्रश्न
An urn contains four tickets marked with numbers 112, 121, 122, 222 and one ticket is drawn at random. Let Ai (i = 1, 2, 3) be the event that ith digit of the number of the ticket drawn is 1. Discuss the independence of the events A1, A2, and A3.
Advertisements
उत्तर
One ticket can be drawn out of 4 tickets in 4C1 = 4 ways.
∴ n(S) = 4
According to given information,
Let A1 be the event that 1st digit of the number of ticket is 1
A2 be the event that 2nd digit of the number of ticket is 1.
A3 be the event that 3rd digit of the number of ticket is 1.
∴ A1 = {112, 121, 122}, A2 = {112}, A3 = {121}
∴ `"P"("A"_1) = ("n"("A"_1))/("n"("S")) = 3/4`,
`"P"("A"_2) = ("n"("A"_2))/("n"("S")) = 1/4`,
`"P"("A"_3) = ("n"("A"_3))/("n"("S")) = 1/4`
`{:("P"("A"_1) "P"("A"_2) = 3/16),("P"("A"_2) "P"("A"_3) = 1/16),("P"("A"_1) "P"("A"_3) = 3/16):}}` ...(i)
A1 ∩ A2 = {112}, A2 ∩ A3 = Φ, A1 ∩ A3 = {121}
`{:("P"("A"_1 ∩ "A"_2) = ("n"("A"_1 ∩ "A"_2))/("n"("S")) = 1/4),("P"("A"_2 ∩ "A"_3) = 0),("P"("A"_1 ∩ "A"_3) = 1/4):}}` ...(ii)
∴ From (i) and (ii),
`{:("P"("A"_1)*"P"("A"_2) ≠ "P"("A"_1 ∩ "A"_2)),("P"("A"_2)*"P"("A"_3) ≠ "P"("A"_2 ∩ "A"_3)),("P"("A"_1)*"P"("A"_3) ≠ "P"("A"_1 ∩ "A"_3)):}}` ...(iii)
∴ A1, A2, A3 are not pairwise independent
For mutual independent of events A1, A2, A3 We require to have
P(A1 ∩ A2 ∩ A3) = P(A1) P(A2) P(A3)
and P(A1) P(A2) = P(A1 ∩ A2),
P(A2) P(A3) = P(A2 ∩ A3),
P(A1) P(A3) = P(A1 ∩ A3)
∴ From (iii),
A1, A2, A3 are not mutually independent.
APPEARS IN
संबंधित प्रश्न
Probability of solving specific problem independently by A and B are `1/2` and `1/3` respectively. If both try to solve the problem independently, find the probability that
- the problem is solved
- exactly one of them solves the problem.
A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is solved?
A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is not solved
One-shot is fired from each of the three guns. Let A, B, and C denote the events that the target is hit by the first, second and third guns respectively. assuming that A, B, and C are independent events and that P(A) = 0.5, P(B) = 0.6, and P(C) = 0.8, then find the probability that at least one hit is registered.
The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?
A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2` and `5/8`. Find the probability that the target
- is hit exactly by one of them
- is not hit by any one of them
- is hit
- is exactly hit by two of them
A bag contains 3 yellow and 5 brown balls. Another bag contains 4 yellow and 6 brown balls. If one ball is drawn from each bag, what is the probability that, both the balls are of the same color?
Solve the following:
A and B throw a die alternatively till one of them gets a 3 and wins the game. Find the respective probabilities of winning. (Assuming A begins the game)
10% of the bulbs produced in a factory are of red colour and 2% are red and defective. If one bulb is picked up at random, determine the probability of its being defective if it is red.
Two dice are thrown together. Let A be the event ‘getting 6 on the first die’ and B be the event ‘getting 2 on the second die’. Are the events A and B independent?
If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.
Two dice are thrown together and the total score is noted. The events E, F and G are ‘a total of 4’, ‘a total of 9 or more’, and ‘a total divisible by 5’, respectively. Calculate P(E), P(F) and P(G) and decide which pairs of events, if any, are independent.
A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A'"/"B'")`
Let E1 and E2 be two independent events such that P(E1) = P1 and P(E2) = P2. Describe in words of the events whose probabilities are: 1 – (1 – P1)(1 – P2)
A and B are events such that P(A) = 0.4, P(B) = 0.3 and P(A ∪ B) = 0.5. Then P(B′ ∩ A) equals ______.
One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?
E : ‘the card drawn is a spade’
F : ‘the card drawn is an ace’
Let EC denote the complement of an event E. Let E1, E2 and E3 be any pairwise independent events with P(E1) > 0 and P(E1 ∩ E2 ∩ E3) = 0. Then `"P"(("E"_2^"C" ∩ "E"_3^"C")/"E"_1)` is equal to ______.
Which is the equivalent and most commonly used test for independent events E and F?
A die is thrown once. If E is the event that the number appearing is a multiple of 3, which set represents E?
Given \[P(E\cap F)=\frac{1}{6}\] and \[P(E)P(F)=\frac{1}{6}\], what is the relation between E and F?
